tom and latoya take turns cleaning the kitchen after dinner. When tom cleans, he finishes the job in 24min. Latoya cleans the kitchen in 30 min. how long would it take the two of them, working, together, to clean the kitchen?
Answer by jorel1380(3719)

t/24+t/30=1
5t+4t=120
t=120/9=40/3 minutes..